Carbex, Inc., produces cutlery sets out of high-quality wood and steel. The company makes a Standard set and a Deluxe set and sells them to retail department stores throughout the country. The Standard set sells for $92, and the Deluxe set sells for $107. The variable expenses associated with each set are given below.

Standard Deluxe
Variable production costs $ 31.00 $ 46.00
Sales commissions (31% of sales price) $ 28.52 $ 33.17

The company’s fixed expenses each month are:

Advertising $ 121,000
Depreciation $ 26,500
Administrative $ 71,000

Mary Parsons, the financial vice president, watches sales commissions carefully and has noted that they have risen steadily over the last year. For this reason, she was shocked to find that even though sales have increased, profits for the current month—May—are down substantially from April. Sales, in sets, for the last two months are given below:

Standard Deluxe Total
April 5,600 3,600 9,200
May 2,600 6,600 9,200

Required:

1-a. Prepare contribution format income statements for April.

1-b. Prepare contribution format income statements for May.

3-a. Compute the break-even point in dollar sales for April.

3-b. Would the break-even point in May be higher or lower than the break-even point in April?
Answer #1

Carbex, Inc.

1a/b) Contribution income statement: ($)

April May
A Sales: $ 900,400 $ 945,400
Standard

(5600×92)

= 515,200

(2600×92)

= 239,200

Deluxe

(3600×107)

= 385,200

(6600×107)

= 706,200

B Variable cost: $ 618,324 $ 677,274
Standard

(5600×59.52)

= 3,33,312

(2600×59.52)

= 154,752

Deluxe

(3600×79.17)

= 285,012

(6600×79.17)

= 522,522

C Contribution (A-B) $ 282,076 $ 268,126
D Fixed cost (121,000+26,500+71,000) 218,500 218,500
E Income (C-D) $ 63,576 $ 49,626

3 a/b) Break even Analysis:

Sales mix:

April May
Standard 5600/9200 = 60.87% 2600/9200 = 28.26%
Deluxe 3600/9200 = 39.13% 6600/9200 = 71.74%
100% 100%

Weighted average contribution per set:

Contribution per unit × Sales mix April May
Standard (92-31-28.52 = 32.48)

32.48 × 60.87%

= 19.77

32.48 × 28.26%

= 9.18

Deluxe (107-46-33.17 = 27.83)

27.83 × 39.13%

= 10.89

27.83 × 71.74%

= 19.97

$ 30.66 $ 29.15

Break even point:

April May
Fixed cost/weighted average contribution per unit

218,500/30.66

= 7,126.55 sets

218,500/29.15

= 7,495.71 sets

Break even point in dollars:

BEP units × Sales mix × Selling price April May
Standard

(7126.55×60.87%×92)

= $ 399,090

(7495.71×28.26%×92)

= $ 194,882

Deluxe

(7126.55×39.13%×107)

= $ 298,382

(7495.71×71.74%×107)

= $ 575,384

$ 697,472 $ 770,266

Break even point in May is higher than the Break even point in April.